# VIPER: Process-aware Evaluation for Generative Video Reasoning

### v1

@@ 6.2 Generative Video Reasoning
As video generation models evolve, their potential for completing reasoning tasks has gradually been uncovered. Research has shown that video generation models can generalize to many common visual tasks that they have not encountered during training, using a sequence of frames to represent the solving process, which is termed Chain-of-Frames (CoF) reasoning Wiedemer et al. (2025). Subsequent work has focused on evaluating and optimizing video models for generative video reasoning tasks. Some studies have designed empirical studies to showcase the reasoning capabilities of video models Guo et al. (2025b); Tong et al. (2025). Additionally, many efforts have proposed benchmarks covering various evaluation dimensions by collecting existing video data or manually constructing examples Chen et al. (2025a); Deng (2025); Zhou et al. (2025); Liu et al. (2025). Furthermore, some studies have made initial explorations into improving models’ reasoning abilities through prompt optimization and supervised fine-tuning using correct video examples, yielding some promising results Chen et al. (2025a); Jialong Wu (2025).
@@ 7 Conclusion
Appendix A Task Introduction
A.1 Temporal Reasoning
@@
Figure 6: Example image for Object Moving.
Object Moving. Example Prompt A clean white background featuring a blue triangle defined by a solid outline at top, and a target destination marked by a matching blue triangle with a dotted outline. The solid-outlined object travels in down, ending perfectly overlapping and aligning with the dotted target outline. The object moves by smoothly rotating clockwise while translating. High contrast, vector graphic style, locked camera, no zoom, consistent scale.
@@
Figure 7: Example image for World Memory.
World Memory. Example Prompt An image. Execute a slow, centered zoom-in to 2x magnification, hold the frame static, and then smoothly zoom out to the original field of view. The visual elements remain static and unchanged throughout the entire process. No pan, no dolly, no perspective shifts.
A.2 Structural Reasoning
@@
Figure 8: Example image for Sudoku.
Sudoku. Example Prompt A clean, high-contrast image of a partially filled 9x9 Sudoku grid on a solid white background. Sovle the Sudoku, with the missing numbers (1-9) appearing in the correct empty cells. Top-down orthographic view, locked camera, no zoom, no pan, no perspective shifts, consistent scale.
@@
This task requires the model to interpret a midgame Tic-Tac-Toe board and place in an appropriate empty cell according to the game state. Images are generated programmatically using Python from valid valid Tic-Tac-Toe configurations.
Figure 9: Example image for Tic Tac Toe.
Tic Tac Toe. Example Prompt A static, top-down view of a minimalist 3x3 Tic-Tac-Toe grid on a clean background. The board is partially filled with existing ’X’ and ’O’ markers in a midgame state. A single ’X’ is drawn into a strategic empty square, executing a logical move to block an opponent or complete a straight line of three. The grid lines and existing marks remain undisturbed. High contrast, vector graphic style, locked camera, no zoom, no dolly, no perspective shifts.
@@
Figure 10: Example image for Checkmate-in-One.
Checkmate-in-One. Example Prompt A static, top-down 2D view of a digital chess board. A single White piece glides smoothly for exactly one step to deliver checkmate to the Black King. Static camera, no pan, no zoom, no dolly, no perspective shifts.
@@
Figure 11: Example image for Maze.
Maze. Example Prompt A top-down 2D maze. The black lines represent walls. the green square marks the starting point, and the red square marks the ending point. Draw a solid red line animates smoothly from the green square navigating the correct path through the maze structure without intersecting the walls, finally reaching the red square. Static camera, no pan, no zoom, no dolly, no perspective shifts.
A.3 Symbolic Reasoning
@@
Figure 12: Example image for Multimodal.
Multimodal. Example Prompt A static shot of a clean, well-lit whiteboard displaying the text or diagram for a multimodal problem: One slice of a circular cake is $15\%$ of the whole cake. What is the size of the angle marked with the question mark?. The solution appears sequentially on the empty space below the problem, writing out the derivation process and concluding with the final answer clearly boxed. Static camera, no pan, no zoom, no perspective shifts, consistent scale.
@@
Figure 13: Example image for Math.
Math. Example Prompt A static shot of a clean, well-lit whiteboard displaying the text or diagram for a math problem: Dennis uses 1 pound of butter for every dozen croissants that he makes. He needs to make 6 dozen croissants. The grocery store currently has a promotion for buy one pound of butter get one half off. If the butter costs \$4.00 a pound, how much will it cost him to purchase 6 pounds of butter?. The solution appears sequentially on the empty space below the problem, writing out the derivation process and concluding with the final answer clearly boxed. Static camera, no pan, no zoom, no perspective shifts, consistent scale.
@@
Figure 14: Example image for Knowledge.
Knowledge. Example Prompt A static shot of a clean, well-lit whiteboard displaying the text or diagram for a knowledge problem: Which of the following (effective) particles is not associated with a spontaneously-broken symmetry? A. Phonon B. Skyrmion C. Pion D. Magnon.. The solution appears sequentially on the empty space below the problem, writing out the derivation process and concluding with the final answer clearly boxed. Static camera, no pan, no zoom, no perspective shifts, consistent scale.
A.4 Spatial Reasoning
@@
Figure 15: Example image for Block Rotate.
Block Rotate. Example Prompt A structure composed of 9 cubes viewed from 316° azimuth, 40° elevation. The camera executes a smooth horizontal orbit around the structure, transitioning to a 136° azimuth which is a 180-degree rotation while maintaining a constant distance from the object. Smooth orbital movement, constant radius, no zoom, no vertical tilt, consistent scale.
Dice Rollong.
@@
Figure 16: Example image for Dice Rollong.
Dice Rollong. Example Prompt A standard six-sided die (faces arranged 1 opposite 6, 2 opposite 5, 3 opposite 4) resting on a grid-marked surface. A thick black path follows the grid lines, originating from the die’s position and terminating at a solid black circular target. The die travels along the black path by executing precise 90-degree flips, pivoting on the edge in contact with the table for each step. The motion continues sequentially along the line until the die lands perfectly covering the black circular target. Static camera, high-angle view, sharp focus, no zoom, no pan, no perspective shifts.
@@
Figure 17: Example image for Image Restore.
Image Restore. Example Prompt A clean white background. A set of shuffled image sub-tiles are arranged at the bottom of the frame, situated below a centered target grid of dashed outlines. The tiles translate smoothly from their starting positions into the matching dashed frames to fully reconstruct the original image. Flat 2D style, locked camera, no zoom, no pan, no perspective shifts.
A.5 Physics Reasoning
@@
Figure 18: Example image for Experiment.
Experiment. Example Prompt A person sits on a chair and keeps spinning. Initially, the arms are rest by his sides, then gradually rise and extend horizontally. Static camera, no pan, no zoom, no dolly.
@@
Figure 19: Example image for Game.
Game. Example Prompt A mobile video game where a candy is fixed by several ropes, each rope connecting one end to the candy and the other end to the wall in the background. A cartoon character is at the bottom of the screen. The candy will be eaten if it moves near the character. Cut the two ropes at the bottom. Static camera, no pan, no zoom, no dolly.
A.6 Planning Reasoning
@@
This task requires the model to render an object manipulation sequence where a robot arm grasps a target item and relocates it to a specified position on the table. Images are collected from RoboBench (Luo et al., 2025).
Figure 20: Example image for Object Manipulation.
Object Manipulation. Example Prompt A lab workbench scene with a wooden table in the center. On the right side of the table there is a gray doll/plush-like object, with other small items nearby. A robot arm is positioned to the right of the table. The robot arm grasps the gray doll and places it down on the left side of the table. Static camera, no pan, no zoom, no dolly.
@@
Figure 21: Example image for Navigation.
Navigation. Example Prompt A 2D map with a top-down view. Draw a red path on the map to navigate from Air, passing through Corkscrew and finally reaching Oblivion. The red path should be continuously drawn as a smooth line. Static camera, no pan, no zoom, no doll.
Appendix B Rubric Examples

### v2

@@ 6.2 Generative Video Reasoning
As video generation models evolve, their potential for completing reasoning tasks has gradually been uncovered. Research has shown that video generation models can generalize to many common visual tasks that they have not encountered during training, using a sequence of frames to represent the solving process, which is termed Chain-of-Frames (CoF) reasoning Wiedemer et al. (2025). Subsequent work has focused on evaluating and optimizing video models for generative video reasoning tasks. Some studies have designed empirical studies to showcase the reasoning capabilities of video models Guo et al. (2025b); Tong et al. (2025). Additionally, many efforts have proposed benchmarks covering various evaluation dimensions by collecting existing video data or manually constructing examples Chen et al. (2025a); Deng (2025); Zhou et al. (2025); Liu et al. (2025); Luo et al. (2025a). Furthermore, some studies have made initial explorations into improving models’ reasoning abilities through prompt optimization and supervised fine-tuning using correct video examples, yielding some promising results Chen et al. (2025a); Wu et al. (2025).
@@ 7 Conclusion
Appendix A Benchmark Statistics
As illustrated in Table 6, VIPER comprises a total of 309 samples. We curate most tasks to contain 10-25 samples, ensuring a balanced distribution to assess model performance across diverse scenarios. The Structural category accounts for the largest portion, followed by the Symbolic and Physics domains.
Domain Task Count Total
Physics Experiment 18 32
Game 14
Planning Navigation 25 44
Object Manipulation 19
Spatial Block Rotate 25 60
Dice Rolling 20
Image Restore 15
Structural Chess 20 70
Maze 20
Sudoku 20
Tic-Tac-Toe 10
Symbolic Knowledge 20 60
Math 20
Multimodal 20
Temporal Object Move 25 43
World Memory 18
Sum 309
Table 6: Detailed statistics of VIPER.
Appendix B Task Introduction
B.1 Temporal Reasoning
@@
Figure 6: Example image of task Object Moving.
Task prompt for Object Moving. A clean white background featuring a blue triangle defined by a solid outline at top, and a target destination marked by a matching blue triangle with a dotted outline. The solid-outlined object travels in down, ending perfectly overlapping and aligning with the dotted target outline. The object moves by smoothly rotating clockwise while translating. High contrast, vector graphic style, locked camera, no zoom, consistent scale.
@@
Figure 7: Example image of task World Memory.
Task prompt for World Memory. An image. Execute a slow, centered zoom-in to 2x magnification, hold the frame static, and then smoothly zoom out to the original field of view. The visual elements remain static and unchanged throughout the entire process. No pan, no dolly, no perspective shifts.
B.2 Structural Reasoning
@@
Figure 8: Example image of task Sudoku.
Task prompt for Sudoku. A clean, high-contrast image of a partially filled 9x9 Sudoku grid on a solid white background. Sovle the Sudoku, with the missing numbers (1-9) appearing in the correct empty cells. Top-down orthographic view, locked camera, no zoom, no pan, no perspective shifts, consistent scale.
@@
This task requires the model to interpret a midgame Tic-Tac-Toe board and place in an appropriate empty cell according to the game state. Images are generated programmatically using Python from valid Tic-Tac-Toe configurations.
Figure 9: Example image of task Tic Tac Toe.
Task prompt for Tic Tac Toe. A static, top-down view of a minimalist 3x3 Tic-Tac-Toe grid on a clean background. The board is partially filled with existing ’X’ and ’O’ markers in a midgame state. A single ’X’ is drawn into a strategic empty square, executing a logical move to block an opponent or complete a straight line of three. The grid lines and existing marks remain undisturbed. High contrast, vector graphic style, locked camera, no zoom, no dolly, no perspective shifts.
@@
Figure 10: Example image of task Checkmate-in-One.
Task prompt for Checkmate-in-One. A static, top-down 2D view of a digital chess board. A single White piece glides smoothly for exactly one step to deliver checkmate to the Black King. Static camera, no pan, no zoom, no dolly, no perspective shifts.
@@
Figure 11: Example image of task Maze.
Task prompt for Maze. A top-down 2D maze. The black lines represent walls. the green square marks the starting point, and the red square marks the ending point. Draw a solid red line animates smoothly from the green square navigating the correct path through the maze structure without intersecting the walls, finally reaching the red square. Static camera, no pan, no zoom, no dolly, no perspective shifts.
B.3 Symbolic Reasoning
@@
Figure 12: Example image of task Multimodal.
Task prompt for Multimodal. A static shot of a clean, well-lit whiteboard displaying the text or diagram for a multimodal problem: One slice of a circular cake is $15\%$ of the whole cake. What is the size of the angle marked with the question mark?. The solution appears sequentially on the empty space below the problem, writing out the derivation process and concluding with the final answer clearly boxed. Static camera, no pan, no zoom, no perspective shifts, consistent scale.
@@
Figure 13: Example image of task Math.
Task prompt for Math. A static shot of a clean, well-lit whiteboard displaying the text or diagram for a math problem: Dennis uses 1 pound of butter for every dozen croissants that he makes. He needs to make 6 dozen croissants. The grocery store currently has a promotion for buy one pound of butter get one half off. If the butter costs \$4.00 a pound, how much will it cost him to purchase 6 pounds of butter?. The solution appears sequentially on the empty space below the problem, writing out the derivation process and concluding with the final answer clearly boxed. Static camera, no pan, no zoom, no perspective shifts, consistent scale.
@@
Figure 14: Example image of task Knowledge.
Task prompt for Knowledge. A static shot of a clean, well-lit whiteboard displaying the text or diagram for a knowledge problem: Which of the following (effective) particles is not associated with a spontaneously-broken symmetry? A. Phonon B. Skyrmion C. Pion D. Magnon.. The solution appears sequentially on the empty space below the problem, writing out the derivation process and concluding with the final answer clearly boxed. Static camera, no pan, no zoom, no perspective shifts, consistent scale.
B.4 Spatial Reasoning
@@
Figure 15: Example image of task Block Rotate.
Task prompt for Block Rotate. A structure composed of 9 cubes viewed from 316° azimuth, 40° elevation. The camera executes a smooth horizontal orbit around the structure, transitioning to a 136° azimuth which is a 180-degree rotation while maintaining a constant distance from the object. Smooth orbital movement, constant radius, no zoom, no vertical tilt, consistent scale.
Dice Rolling.
@@
Figure 16: Example image of task Dice Rolling.
Task prompt for Dice Rolling. A standard six-sided die (faces arranged 1 opposite 6, 2 opposite 5, 3 opposite 4) resting on a grid-marked surface. A thick black path follows the grid lines, originating from the die’s position and terminating at a solid black circular target. The die travels along the black path by executing precise 90-degree flips, pivoting on the edge in contact with the table for each step. The motion continues sequentially along the line until the die lands perfectly covering the black circular target. Static camera, high-angle view, sharp focus, no zoom, no pan, no perspective shifts.
@@
Figure 17: Example image of task Image Restore.
Task prompt for Image Restore. A clean white background. A set of shuffled image sub-tiles are arranged at the bottom of the frame, situated below a centered target grid of dashed outlines. The tiles translate smoothly from their starting positions into the matching dashed frames to fully reconstruct the original image. Flat 2D style, locked camera, no zoom, no pan, no perspective shifts.
B.5 Physics Reasoning
@@
Figure 18: Example image of task Experiment.
Task prompt for Experiment. A person sits on a chair and keeps spinning. Initially, the arms are rest by his sides, then gradually rise and extend horizontally. Static camera, no pan, no zoom, no dolly.
@@
Figure 19: Example image of task Game.
Task prompt for Game. A mobile video game where a candy is fixed by several ropes, each rope connecting one end to the candy and the other end to the wall in the background. A cartoon character is at the bottom of the screen. The candy will be eaten if it moves near the character. Cut the two ropes at the bottom. Static camera, no pan, no zoom, no dolly.
B.6 Planning Reasoning
@@
This task requires the model to render an object manipulation sequence where a robot arm grasps a target item and relocates it to a specified position on the table. Images are collected from RoboBench (Luo et al., 2025b).
Figure 20: Example image of task Object Manipulation.
Task prompt for Object Manipulation. A lab workbench scene with a wooden table in the center. On the right side of the table there is a gray doll/plush-like object, with other small items nearby. A robot arm is positioned to the right of the table. The robot arm grasps the gray doll and places it down on the left side of the table. Static camera, no pan, no zoom, no dolly.
@@
Figure 21: Example image of task Navigation.
Task prompt for Navigation. A 2D map with a top-down view. Draw a red path on the map to navigate from Air, passing through Corkscrew and finally reaching Oblivion. The red path should be continuously drawn as a smooth line. Static camera, no pan, no zoom, no doll.
Appendix C Rubric Examples